In a survey of men in the United States (ages 20-29), the mean height was 69.6 inches with a standard deviation of 3.0 inches. Find the minimum height in the top 22%.


1
Expert's answer
2021-03-01T07:24:41-0500

"μ = 69.6 \\

σ = 3.0 \\

P(X<x) = P(Z< \frac{x-69.6}{3.0}) = 1 -0.22 = 0.79 \\

P(Z< \frac{x-69.6}{3.0}) = 0.79 \\

\frac{x-69.6}{3.0} = 0.81 \\

x -69.6 = 2.43 \\

x = 72.03"

The minimum height in the top 22% is 72.03 inches.